a certain number is multiplied by 4 and the product is divided by 3. if the result is 20 . what is the number?
Answer by deborabr(181)

the certain number is x
if it is multiplied by 4 we get 4x
if the product is divided by 3 we get 4x/3
and if the result is 20 we get 4x/3=20
4x/3=20 --> we can pass 3 multiplying because it is dividing
4x=20*3
4x=60
x=60/4
x=15
the number is 15